# definition

### v2

@@ D.1 Preliminary lemmas on transition matrices
First, we show that the norm of $(P_{i}-\Xi_{i})^{n}$ decays exponentially fast when $n$ is larger than some constant independent of $N$, as claimed below:
Claim:
There exists a constant $n(\gamma,\lvert\mathbb{S}\rvert,\gamma_{\rho})>0$ which only depends on the parameter $\gamma$, the state space size $\lvert\mathbb{S}\rvert$, and the uniform upper bound on the spectral radius $\gamma_{\rho}$, such that for any $n\geq n(\gamma,\lvert\mathbb{S}\rvert,\gamma_{\rho})$, we have:
$\left\lVert(P_{i}-\Xi_{i})^{n}\right\rVert\leq\left(\frac{\gamma+\gamma_{\rho}}{2}\right)^{n}\quad\forall i=1,2,\dots.$
To prove this claim, we consider two cases:
Case 1:
$\left\lVert(P_{i}-\Xi_{i})^{\lvert\mathbb{S}\rvert}\right\rVert\leq\left(\frac{\gamma+\gamma_{\rho}}{4}\right)^{\lvert\mathbb{S}\rvert}$. In this case, using the sub-multiplicative property of matrix norms, we get:
@@
$\left\lVert(P_{i}-\Xi_{i})^{n}\right\rVert\leq\left\lVert(P_{i}-\Xi_{i})^{\lvert\mathbb{S}\rvert}\right\rVert^{\lfloor\frac{n}{\lvert\mathbb{S}\rvert}\rfloor}\left\lVert P_{i}-\Xi_{i}\right\rVert^{(n~{}\text{mod}~{}\lvert\mathbb{S}\rvert)}\leq\left(\frac{\gamma+\gamma_{\rho}}{4}\right)^{(n-\lvert\mathbb{S}\rvert)}\left(2\lvert\mathbb{S}\rvert\right)^{\lvert\mathbb{S}\rvert}\,,$
where $(n~{}\text{mod}~{}\lvert\mathbb{S}\rvert)$ denotes the remainder of $n$ after dividing $\lvert\mathbb{S}\rvert$; the second inequality is due to $\left\lVert P_{i}-\Xi_{i}\right\rVert\leq\left\lVert P_{i}\right\rVert+\left\lVert\Xi_{i}\right\rVert\leq 2{\lvert\mathbb{S}\rvert}$. Therefore, if $n\geq\lvert\mathbb{S}\rvert\left(3+\log\frac{1}{\gamma}+\log\lvert\mathbb{S}\rvert\right)$, the above inequality implies that $\left\lVert(P_{i}-\Xi_{i})^{n}\right\rVert\leq\left(\frac{\gamma+\gamma_{\rho}}{2}\right)^{n}$.
Case 2:
$\left\lVert(P_{i}-\Xi_{i})^{\lvert\mathbb{S}\rvert}\right\rVert>\left(\frac{\gamma+\gamma_{\rho}}{4}\right)^{\lvert\mathbb{S}\rvert}$. In this case, we have
$\frac{\left\lVert P_{i}-\Xi_{i}\right\rVert^{\lvert\mathbb{S}\rvert}}{\left\lVert\left(P_{i}-\Xi_{i}\right)^{\lvert\mathbb{S}\rvert}\right\rVert}\leq\frac{(2\lvert\mathbb{S}\rvert)^{\lvert\mathbb{S}\rvert}}{\left\lVert\left(P_{i}-\Xi_{i}\right)^{\lvert\mathbb{S}\rvert}\right\rVert}<\left(\frac{8}{\gamma+\gamma_{\rho}}\right)^{\lvert\mathbb{S}\rvert}\,.$
Combining the above inequality with Lemma 8, we get:
$\left\lVert\left(P_{i}-\Xi_{i}\right)^{n}\right\rVert$ $\leq\rho(P_{i}-\Xi_{i})^{n}C_{\lvert\mathbb{S}\rvert}^{\sigma_{\lvert\mathbb{S}\rvert}(n)}\left(\frac{\left\lVert P_{i}-\Xi_{i}\right\rVert^{\lvert\mathbb{S}\rvert}}{\left\lVert\left(P_{i}-\Xi_{i}\right)^{\lvert\mathbb{S}\rvert}\right\rVert}\right)^{\nu_{\lvert\mathbb{S}\rvert}(n)}$
$\leq\gamma_{\rho}^{n}C_{\lvert\mathbb{S}\rvert}^{\sigma_{\lvert\mathbb{S}\rvert}(n)}\left(\frac{8}{\gamma+\gamma_{\rho}}\right)^{\lvert\mathbb{S}\rvert\nu_{\lvert\mathbb{S}\rvert}(n)},$ (33)
where $C_{\lvert\mathbb{S}\rvert}$, $\sigma_{\lvert\mathbb{S}\rvert}(n)$ and $\nu_{\lvert\mathbb{S}\rvert}(n)$ are given in (8). Because we have $\gamma_{\rho}<\frac{\gamma+\gamma_{\rho}}{2}$, and $C_{\lvert\mathbb{S}\rvert}^{\sigma_{\lvert\mathbb{S}\rvert}(n)}\left(\frac{8}{\gamma+\gamma_{\rho}}\right)^{\lvert\mathbb{S}\rvert\nu_{\lvert\mathbb{S}\rvert}(n)}$ grows with $n$ at a sub-exponential rate, there exists $n_{0}(\gamma,\lvert\mathbb{S}\rvert,\gamma_{\rho})>0$ such that for each $n\geq n_{0}(\gamma,\lvert\mathbb{S}\rvert,\gamma_{\rho})$, we have
$\gamma_{\rho}^{n}C_{\lvert\mathbb{S}\rvert}^{\sigma_{\lvert\mathbb{S}\rvert}(n)}\left(\frac{8}{\gamma+\gamma_{\rho}}\right)^{\lvert\mathbb{S}\rvert\nu_{\lvert\mathbb{S}\rvert}(n)}\leq\left(\frac{\gamma+\gamma_{\rho}}{2}\right)^{n}.$
Putting together the two cases and choosing $n(\gamma,\lvert\mathbb{S}\rvert,\gamma_{\rho})=\max\left\{n_{0}(\gamma,\lvert\mathbb{S}\rvert,\gamma_{\rho}),\lvert\mathbb{S}\rvert\left(3+\log\frac{1}{\gamma}+\log\lvert\mathbb{S}\rvert\right)\right\}$, we finish the proof of the claim.
Using the above claim, we get:
@@
$\sum_{n=0}^{\infty}\sup_{i\in[N]}\frac{\left\lVert(P_{i}-\Xi_{i})^{n}\right\rVert}{\gamma^{n}}$ $\leq\sum_{n=0}^{n(\gamma,\lvert\mathbb{S}\rvert,\gamma_{\rho})-1}\sup_{i\in[N]}\frac{\left\lVert(P_{i}-\Xi_{i})^{n}\right\rVert}{\gamma^{n}}+\sum_{n=n(\gamma,\lvert\mathbb{S}\rvert,\gamma_{\rho})}^{\infty}\left(\frac{\gamma+\gamma_{\rho}}{2\gamma}\right)^{n}$ (34)
@@
$\leq\sum_{n=0}^{n(\gamma,\lvert\mathbb{S}\rvert,\gamma_{\rho})-1}\frac{(2\lvert\mathbb{S}\rvert)^{n}}{\gamma^{n}}+\sum_{n=n(\gamma,\lvert\mathbb{S}\rvert,\gamma_{\rho})}^{\infty}\left(\frac{\gamma+\gamma_{\rho}}{2\gamma}\right)^{n},$ (35)
where the infinite sum on the right-hand side of (34) is finite because $\gamma+\gamma_{\rho}<2\gamma$; to get (35), we used the argument that $\left\lVert(P_{i}-\Xi_{i})^{n}\right\rVert\leq\left\lVert P_{i}-\Xi_{i}\right\rVert^{n}\leq\big{(}\left\lVert P_{i}\right\rVert+\left\lVert\Xi_{i}\right\rVert\big{)}^{n}\leq(2\lvert\mathbb{S}\rvert)^{n}$. Since the final expression in (35) is independent of $N$, taking it to be $C_{\gamma}$ finishes the proof of the lemma. ∎
D.2 Proof of Lemma 3
@@
In the proof of Lemma 3, we will frequently use the following form of $h({\bm{X}},D)$:
@@
where $\Xi_{i}$ is the matrix whose each row is the optimal stationary distribution of the $i$-th arm, $\mu^{*}_{i}$. The equation (36) is equivalent to (18) because $(v_{1}-v_{2})P_{i}^{\ell}=(v_{1}-v_{2})(P_{i}-\Xi_{i})^{\ell}$ for any $i\in[N]$, $\ell\geq 0$, and row vectors $v_{1},v_{2}\in\Delta(\mathbb{S})$. We will also use the equivalent version of 1 proved in Lemma 6, i.e., the spectral radius of the matrix $P_{i}-\Xi_{i}$ is upper bounded by $\gamma_{\rho}$ for any $i=1,2,3,\dots$.
@@
We organize the proof in three parts: we first show the finiteness of the subset Lyapunov function $h(\bm{x},D)$; then, we prove the Lipschitz continuity of $h(\bm{x},D)$ with respect to $D$ (19); finally, we prove the drift condition for $h(\bm{x},D)$ stated in (20).
@@
$\leq\sum_{i\in D}\sum_{\ell=0}^{\infty}\left\lVert x_{i}-\mu_{i}^{*}\right\rVert\frac{\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert}{\gamma^{\ell}}\left\lVert g_{i}\right\rVert,$ (38)
By Lemma 7, $\sum_{\ell=0}^{\infty}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert/\gamma^{\ell}$ is finite, so the expression in (38) is also finite. Taking maximum over $g\in\mathcal{G}$, because $\mathcal{G}$ is a finite set, we have
@@
$\max_{g\in\mathcal{G}}\sup_{\ell\in\mathbb{N}}\left\lvert\sum_{i\in D/D^{\prime}}\left\langle(x_{i}-\mu_{i}^{*})P_{i}^{\ell}/\gamma^{\ell},g_{i}\right\rangle\right\rvert$ $\leq\max_{g\in\mathcal{G}}\sum_{i\in D/D^{\prime}}\left\lVert x_{i}-\mu_{i}^{*}\right\rVert\left\lVert g_{i}\right\rVert\sum_{\ell=0}^{\infty}\frac{\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert}{\gamma^{\ell}}$
$\leq 2\left\lvert D/D^{\prime}\right\rvert\max\left\{c_{\max},r_{\max}\right\}\lvert\mathbb{S}\rvert^{1/2}C_{\gamma},$ (40)
where in the last inequality, we have utilized the facts that $\left\lVert g_{i}\right\rVert\leq\max\left\{c_{\max},r_{\max}\right\}\lvert\mathbb{S}\rvert^{1/2}$, $\left\lVert x_{i}-\mu_{i}^{*}\right\rVert\leq 2$, and that $\sum_{\ell=0}^{\infty}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert/\gamma^{\ell}\leq C_{\gamma}$ for some constant $C_{\gamma}>0$. Therefore, $h(\bm{x},D)$ is Lipschitz continuous in $D$ with the Lipschitz constant $L_{h}=2\max\left\{c_{\max},r_{\max}\right\}\lvert\mathbb{S}\rvert^{1/2}C_{\gamma}$.
@@
$\leq\gamma h({\bm{X}}_{t}).$
@@
$\leq\left\lvert\max_{g\in\mathcal{G}}\sup_{\ell\in\mathbb{N}}\left(\left\lvert\sum_{i\in D}\left\langle(X_{i,t+1}-\mu_{i}^{*})P_{i}^{\ell}/\gamma^{\ell},g_{i}\right\rangle\right\rvert-\left\lvert\sum_{i\in D}\left\langle(X_{i,t}P_{i}-\mu_{i}^{*})P_{i}^{\ell}/\gamma^{\ell},g_{i}\right\rangle\right\rvert\right)\right\rvert$
@@
$=\max_{g\in\mathcal{G}}\sup_{\ell\in\mathbb{N}}\left\lvert\sum_{i\in D}\sum_{s}\sum_{s^{\prime}\in\mathbb{S}}\epsilon_{i,t}(s)g_{i}(s^{\prime})\left(\frac{P_{i}-\Xi_{i}}{\gamma}\right)^{\ell}(s,s^{\prime})\right\rvert$
$\leq\sum_{g\in\mathcal{G}}\sum_{\ell=0}^{\infty}\sum_{s\in\mathbb{S}}\left\lvert\sum_{i\in D}\epsilon_{i,t}(s)\sum_{s^{\prime}\in\mathbb{S}}\left(g_{i}(s^{\prime})\left(\frac{P_{i}-\Xi_{i}}{\gamma}\right)^{\ell}(s,s^{\prime})\right)\right\rvert$
Let $q_{g,i}^{\ell}(s)=\sum_{s^{\prime}}g_{i}(s^{\prime})\left(\frac{P_{i}-\Xi_{i}}{\gamma}\right)^{\ell}(s,s^{\prime})$ for $\ell\geq 0$, $g\in\mathcal{G}$, $i\in D$ and $s\in\mathbb{S}$. Then
$\left\lvert h({\bm{X}}_{t+1})-h({\bm{X}}_{t}P)\right\rvert\leq\sum_{g\in\mathcal{G}}\sum_{\ell=0}^{\infty}\sum_{s\in\mathbb{S}}\left\lvert\sum_{i\in D}\epsilon_{i,t}(s)q_{g,i}^{\ell}(s)\right\rvert.$ (42)
@@
We can bound the conditional expectation of $\left\lvert\sum_{i\in D}\epsilon_{i,t}(s)q_{g,i}^{\ell}(s)\right\rvert$ given ${\bm{X}}_{t}$ as follows: for any $\ell\geq 0$, $g\in\mathcal{G}$, $i\in D$, we have
@@
$\mathbb{E}\left[\left\lvert\sum_{i\in D}\epsilon_{i,t}(s)q_{g,i}^{\ell}(s)\right\rvert\;\middle|\;{\bm{X}}_{t}\right]$ $=\mathbb{E}\left[\left\lvert\sum_{i\in D}\epsilon_{i,t}(s)q_{g,i}^{\ell}(s)\right\rvert\;\middle|\;{\bm{X}}_{t}\right]$
$\leq\sqrt{\mathbb{E}\left[\left(\sum_{i\in D}\epsilon_{i,t}(s)q_{g,i}^{\ell}(s)\right)^{2}\;\middle|\;{\bm{X}}_{t}\right]}$
$=\sqrt{\sum_{i\in D}\left(q_{g,i}^{\ell}(s)\right)^{2}\mathbb{E}\left[\epsilon_{i,t}(s)^{2}\;\middle|\;{\bm{X}}_{t}\right]}$ (43)
$\leq\sqrt{N}\max_{i\in D}\left\lvert q_{g,i}^{\ell}(s)\right\rvert,$ (44)
where (43) uses the fact that $\epsilon_{i,t}$ are independent across $i\in D$; (44) is because $\left\lvert\epsilon_{i,t}(s)\right\rvert\leq 1$ for $i\in D$ and $s\in\mathbb{S}$. Thus, we can bound $\mathbb{E}\left[h({\bm{X}}_{t+1})-h({\bm{X}}_{t}P)\;\middle|\;{\bm{X}}_{t}\right]$ as:
@@
$\mathbb{E}\left[\left\lvert h({\bm{X}}_{t+1})-h({\bm{X}}_{t}P)\right\rvert\;\middle|\;{\bm{X}}_{t}\right]$ $\leq\sqrt{N}\sum_{\ell=0}^{\infty}\sum_{s\in\mathbb{S}}\sum_{g\in\mathcal{G}}\max_{i\in D}\left\lvert q_{g,i}^{\ell}(s)\right\rvert.$ (45)
To bound the term $\sum_{s}\sum_{g}\max_{i\in D}\left\lvert q_{g,i}^{\ell}(s)\right\rvert$ in (45), we calculate that
$\left\lvert q_{g,i}^{\ell}(s)\right\rvert$ $\leq\left(\max_{s^{\prime}\in\mathbb{S}}\left\lvert g_{i}(s^{\prime})\right\rvert\right)\left(\sum_{s^{\prime}\in\mathbb{S}}\left\lvert\left(\frac{P_{i}-\Xi_{i}}{\gamma}\right)^{\ell}(s,s^{\prime})\right\rvert\right)$
$\leq\left\lVert g_{i}\right\rVert_{\infty}\frac{\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}}{\gamma^{\ell}}\,.$
Consequently, summing over $s\in\mathbb{S}$ and $g\in\mathcal{G}$,
$\sum_{s\in\mathbb{S}}\sum_{g\in\mathcal{G}}\max_{i\in D}\left\lvert q_{g,i}^{\ell}(s)\right\rvert$ $\leq\lvert\mathbb{S}\rvert\sum_{g\in\mathcal{G}}\left(\max_{i\in[N]}\left\lVert g_{i}\right\rVert_{\infty}\right)\left(\sup_{i\in[N]}\frac{\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}}{\gamma^{\ell}}\right)$
$\leq\lvert\mathbb{S}\rvert(Kc_{\max}+r_{\max})\left(\sup_{i\in[N]}\frac{\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}}{\gamma^{\ell}}\right)\,.$ (46)
Plugging the bound in (46) back to (45), we get
@@
$\mathbb{E}\left[\left\lvert h({\bm{X}}_{t+1})-h({\bm{X}}_{t}P)\right\rvert\;\middle|\;{\bm{X}}_{t}\right]$ $\leq\sqrt{N}\lvert\mathbb{S}\rvert(Kc_{\max}+r_{\max})\sum_{\ell=0}^{\infty}\sup_{i\in[N]}\frac{\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}}{\gamma^{\ell}}$
@@
$\leq\sqrt{N}\lvert\mathbb{S}\rvert^{3/2}(Kc_{\max}+r_{\max})C_{\gamma}\,,$ (47)
@@
where the inequality in (47) is due to the fact that for any $\lvert\mathbb{S}\rvert$-by-$\lvert\mathbb{S}\rvert$ matrix $A$, $\left\lVert A\right\rVert_{\infty}\leq\lvert\mathbb{S}\rvert^{1/2}\left\lVert A\right\rVert_{2}$, and the upper bound $\sum_{\ell=0}^{\infty}\sup_{i}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{2}/\gamma^{\ell}\leq C_{\gamma}$ proved in Lemma 7.
@@
$\leq\sqrt{N}\lvert\mathbb{S}\rvert^{3/2}(Kc_{\max}+r_{\max})C_{\gamma}.$
Therefore, $\mathbb{E}\left[\left(h({\bm{X}}_{t+1})-\gamma h({\bm{X}}_{t})\right)^{+}\;\middle|\;{\bm{X}}_{t}\right]\leq C_{h}\sqrt{N}$ with $C_{h}=\lvert\mathbb{S}\rvert^{3/2}(Kc_{\max}+r_{\max})C_{\gamma}$. ∎
D.3 Properties of $h_{\textnormal{ID}}(\cdot,\cdot)$

### v3

@@ E.1 Lemma for utilizing mixing time bound of transition matrices
For each $\ell\in\mathbb{N}$, let $\ell_{0}=\ell-\tau\lfloor\ell/\tau\rfloor$. By properties of the operator norm $\left\lVert\cdot\right\rVert_{\infty}$,
$\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}\leq\left\lVert(P_{i}-\Xi_{i})^{\ell_{0}}\right\rVert_{\infty}\left\lVert(P_{i}-\Xi_{i})^{\tau}\right\rVert_{\infty}^{\lfloor\ell/\tau\rfloor}.$ (30)
Because each row of $\Xi_{i}$ is $\mu_{i}^{*}$, it is easy to verify that $(P_{i}-\Xi_{i})^{\ell^{\prime}}=P_{i}^{\ell^{\prime}}-\Xi_{i}$ for any $\ell^{\prime}\in\mathbb{N}$. Consequently,
$\left\lVert(P_{i}-\Xi_{i})^{\ell_{0}}\right\rVert_{\infty}$ $=\big{\lVert}P_{i}^{\ell_{0}}-\Xi_{i}\big{\rVert}_{\infty}$
$=\max_{s\in\mathbb{S}}\sum_{s^{\prime}\in\mathbb{S}}\left\lvert P_{i}^{\ell_{0}}(s,s^{\prime})-\mu_{i}^{*}(s^{\prime})\right\rvert$
$\leq\max_{s\in\mathbb{S}}\sum_{s^{\prime}\in\mathbb{S}}\big{(}P_{i}^{\ell_{0}}(s,s^{\prime})+\mu_{i}^{*}(s^{\prime})\big{)}$
$=2.$
Moreover, by 1, it is not hard to see that
$\max_{s\in\mathbb{S}}\left\lVert P_{i}^{\tau}(s,\cdot)-\mu_{i}^{*}(\cdot)\right\rVert_{1}\leq\frac{1}{e},$
so we have
$\left\lVert(P_{i}-\Xi_{i})^{\tau}\right\rVert_{\infty}$ $=\max_{s\in\mathbb{S}}\sum_{s^{\prime}\in\mathbb{S}}\left\lvert P_{i}^{\tau}(s,s^{\prime})-\mu_{i}^{*}(s^{\prime})\right\rvert$
$=\max_{s\in\mathbb{S}}\left\lVert P_{i}^{\tau}(s,\cdot)-\mu_{i}^{*}(\cdot)\right\rVert_{1}$
$\leq\frac{1}{e}.$
@@
Substituting the bounds on $\left\lVert(P_{i}-\Xi_{i})^{\tau}\right\rVert_{\infty}$ and $\left\lVert(P_{i}-\Xi_{i})^{\ell_{0}}\right\rVert_{\infty}$ back to (30), we get
@@
$\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}$ $\leq 2\exp\Big{(}-\Big{\lfloor}\frac{\ell}{\tau}\Big{\rfloor}\Big{)}$
$\leq 2\exp\Big{(}-\frac{\ell}{\tau}+1\Big{)}.$ (31)
Therefore,
$\sum_{\ell=0}^{\infty}\max_{i\in[N]}\frac{\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}}{\gamma^{\ell}}$ $\leq\sum_{\ell=0}^{\infty}2\exp\Big{(}-\frac{\ell}{\tau}+1\Big{)}\gamma^{-\ell}$
$=\sum_{\ell=0}^{\infty}2\exp\Big{(}-\frac{\ell}{2\tau}+1\Big{)}$
$=\frac{2e}{1-\exp(-1/(2\tau))}.$ (32)
Observe that $\tau\geq 1$, 0<1/($20<1/(2\tau)\leq 1/$ 2, and $\exp(-a)\leq 1-(1-1/\sqrt{e})a$ for any $20<a\leq 1/$ 2, so
$\frac{2e}{1-\exp(-1/(2\tau))}\leq\frac{4e}{1-1/\sqrt{e}}\tau=C_{\tau}.\qed$
E.2 Proof of Lemma 3
@@
In the proof of Lemma 3, we will frequently use the following equivalent form of $h({\bm{X}},D)$:
@@
where $\Xi_{i}$ is the matrix whose each row is the optimal state distribution of the $i$-th arm, $\mu^{*}_{i}$. The equation (33) is equivalent to (22) because $(v_{1}-v_{2})P_{i}^{\ell}=(v_{1}-v_{2})(P_{i}-\Xi_{i})^{\ell}$ for any $i\in[N]$, $\ell\in\mathbb{N}$, and row vectors $v_{1},v_{2}\in\Delta(\mathbb{S})$.
@@
We organize the proof into three parts: we first show the finiteness of the subset Lyapunov function $h(\bm{x},D)$; then, we prove the Lipschitz continuity of $h(\bm{x},D)$ with respect to $D$ (23); finally, we prove the drift condition for $h(\bm{x},D)$ stated in (24).
@@
$\leq\sum_{i\in D}\sum_{\ell=0}^{\infty}\left\lVert x_{i}-\mu_{i}^{*}\right\rVert_{1}\frac{\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}}{\gamma^{\ell}}\left\lVert g_{i}\right\rVert_{\infty},$ (35)
By Lemma 6, $\sum_{\ell=0}^{\infty}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}/\gamma^{\ell}$ is finite, so the expression in (35) is also finite. Taking maximum over $g\in\mathcal{G}$, because $\mathcal{G}$ is a finite set, we have
@@
$\max_{g\in\mathcal{G}}\sup_{\ell\in\mathbb{N}}\left\lvert\sum_{i\in D/D^{\prime}}\left\langle(x_{i}-\mu_{i}^{*})P_{i}^{\ell}/\gamma^{\ell},g_{i}\right\rangle\right\rvert$ $\leq\max_{g\in\mathcal{G}}\sum_{i\in D/D^{\prime}}\left\lVert x_{i}-\mu_{i}^{*}\right\rVert_{1}\left\lVert g_{i}\right\rVert_{\infty}\sum_{\ell=0}^{\infty}\frac{\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}}{\gamma^{\ell}}$
$\leq 2\left\lvert D/D^{\prime}\right\rvert\max\left\{c_{\max},r_{\max}\right\}C_{\tau},$ (37)
where in the last inequality, we utilize the facts that for any $g\in\mathcal{G}$ and $i\in[N]$, $\left\lVert g_{i}\right\rVert_{\infty}\leq\max\left\{c_{\max},r_{\max}\right\}$, $\left\lVert x_{i}-\mu_{i}^{*}\right\rVert_{1}\leq 2$, and $\sum_{\ell=0}^{\infty}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}/\gamma^{\ell}\leq C_{\tau}$. Therefore, $h(\bm{x},D)$ is Lipschitz continuous in $D$ with the Lipschitz constant $L_{h}=2\max\left\{c_{\max},r_{\max}\right\}C_{\tau}$.
@@
$\leq\gamma h({\bm{X}}_{t}),$ (39)
so $\left(h({\bm{X}}_{t}P)-\gamma h({\bm{X}}_{t})\right)^{+}=0$ with probability $1$.
@@
$\leq\max_{g\in\mathcal{G}}\sup_{\ell\in\mathbb{N}}\left\lvert\bigg{\lvert}\sum_{i\in D}\left\langle(X_{i,t+1}-\mu_{i}^{*})P_{i}^{\ell}/\gamma^{\ell},g_{i}\right\rangle\bigg{\rvert}-\bigg{\lvert}\sum_{i\in D}\left\langle(X_{i,t}P_{i}-\mu_{i}^{*})P_{i}^{\ell}/\gamma^{\ell},g_{i}\right\rangle\bigg{\rvert}\right\rvert$
@@
$=\max_{g\in\mathcal{G}}\sup_{\ell\in\mathbb{N}}\gamma^{-\ell}\left\lvert\sum_{i\in D}\left\langle\epsilon_{i,t},(P_{i}-\Xi_{i})^{\ell}g_{i}\right\rangle\right\rvert.$ (40)
@@
Taking the conditional expectation of $\left\lvert h({\bm{X}}_{t+1})-h({\bm{X}}_{t}P)\right\rvert$ given ${\bm{X}}_{t}$, then (40) implies that
@@
$\mathbb{E}\left[\left\lvert h({\bm{X}}_{t+1})-h({\bm{X}}_{t}P)\right\rvert\;\middle|\;{\bm{X}}_{t}\right]$ $\leq\mathbb{E}\left[\max_{g\in\mathcal{G}}\sup_{\ell\in\mathbb{N}}\gamma^{-\ell}\left\lvert\sum_{i\in D}\left\langle\epsilon_{i,t},(P_{i}-\Xi_{i})^{\ell}g_{i}\right\rangle\right\rvert\;\middle|\;{\bm{X}}_{t}\right]$
@@
$\leq\sum_{g\in\mathcal{G}}\sum_{\ell=0}^{\infty}\gamma^{-\ell}\mathbb{E}\left[\left\lvert\sum_{i\in D}\left\langle\epsilon_{i,t},(P_{i}-\Xi_{i})^{\ell}g_{i}\right\rangle\right\rvert\;\middle|\;{\bm{X}}_{t}\right].$ (41)
We bound the conditional expectation in (41) as follows: for any $\ell\in\mathbb{N}$ and $g\in\mathcal{G}$, we have
@@
$\mathbb{E}\left[\left\lvert\sum_{i\in D}\left\langle\epsilon_{i,t},(P_{i}-\Xi_{i})^{\ell}g_{i}\right\rangle\right\rvert\;\middle|\;{\bm{X}}_{t}\right]$ $\leq\sqrt{\mathbb{E}\left[\bigg{(}\sum_{i\in D}\left\langle\epsilon_{i,t},(P_{i}-\Xi_{i})^{\ell}g_{i}\right\rangle\bigg{)}^{2}\;\middle|\;{\bm{X}}_{t}\right]}$
$=\sqrt{\sum_{i\in D}\mathbb{E}\left[\left\langle\epsilon_{i,t},(P_{i}-\Xi_{i})^{\ell}g_{i}\right\rangle^{2}\;\middle|\;{\bm{X}}_{t}\right]}$ (42)
@@
$\leq\sqrt{\sum_{i\in D}\mathbb{E}\left[\Big{(}\left\lVert\epsilon_{i,t}\right\rVert_{1}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}\left\lVert g_{i}\right\rVert_{\infty}\Big{)}^{2}\;\middle|\;{\bm{X}}_{t}\right]}$
$\leq 2\sqrt{\left\lvert D\right\rvert}\max_{i\in D}\left\lVert g_{i}\right\rVert_{\infty}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty},$ (43)
where (42) uses the facts that $\epsilon_{i,t}$ are independent across $i\in D$ and $\left\langle\epsilon_{i,t},(P_{i}-\Xi_{i})^{\ell}g_{i}\right\rangle$ has zero mean conditioned on ${\bm{X}}_{t}$; (43) is because $\left\lVert\epsilon_{i,t}\right\rVert_{1}=\left\lVert X_{i,t+1}-X_{i,t}P_{i}\right\rVert_{1}\leq 2$ for each $i\in D$.
Substituting the bound in (43) back to (41), we get
$\mathbb{E}\left[\left\lvert h({\bm{X}}_{t+1})-h({\bm{X}}_{t}P)\right\rvert\;\middle|\;{\bm{X}}_{t}\right]$ $\leq 2\sqrt{\left\lvert D\right\rvert}\sum_{\ell=0}^{\infty}\sum_{g\in\mathcal{G}}\gamma^{-\ell}\max_{i\in D}\left\lVert g_{i}\right\rVert_{\infty}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}$
$=2\sqrt{\left\lvert D\right\rvert}\sum_{\ell=0}^{\infty}\sum_{g\in\mathcal{G}}\max_{i\in D}\left\lVert g_{i}\right\rVert_{\infty}\cdot\max_{i\in D}\gamma^{-\ell}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}$
$\leq 2\sqrt{\left\lvert D\right\rvert}(Kc_{\max}+r_{\max})\sum_{\ell=0}^{\infty}\max_{i\in D}\frac{\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}}{\gamma^{\ell}}$
$\leq 2\sqrt{N}(Kc_{\max}+r_{\max})C_{\tau},$ (44)
@@
where the inequality in (44) is due to $\sum_{\ell=0}^{\infty}\max_{i\in D}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}/\gamma^{\ell}\leq C_{\tau}$ (Lemma 6).
@@
$\leq 2\sqrt{N}(Kc_{\max}+r_{\max})C_{\tau}.$
Therefore, $\mathbb{E}\left[\left(h({\bm{X}}_{t+1})-\gamma h({\bm{X}}_{t})\right)^{+}\;\middle|\;{\bm{X}}_{t}\right]\leq C_{h}\sqrt{N}$ with $C_{h}=2(Kc_{\max}+r_{\max})C_{\tau}$. ∎
E.3 Properties of $h_{\textnormal{ID}}(\cdot,\cdot)$